OpenStudy (anonymous):

Not sure what closed meant, thought it was to inform other people I've been helped already :\

OpenStudy (anonymous):

you have to find what 2 times X subtracted by 3 times Y =16 so try a few numbers like 2 for X and 4 for Y

OpenStudy (anonymous):

i hope that helped you

OpenStudy (anonymous):

pretty confused

OpenStudy (anonymous):

yea it is

OpenStudy (anonymous):

have you learned the elimination method yet?

OpenStudy (anonymous):

Unfortunately can't say I have

OpenStudy (anonymous):

well, it's fairly easy and would be really nice for this problem. you down for giving it a go?

OpenStudy (anonymous):

First you have to set up the equations, i think the easiest way is substitution. \[2x-3y=16 \]\[-3x+2y=-19\] You then do 2-3x which gives you -1x after that you move to the next set. -3y+2y which gives you -1y. Then you move past the equal signs and do 16-19 which gives you -3. at this point you should have \[-1x-1y=-3\] now you want an equation for substitution. so you get x by itself. add 1x to both sides. so you are left with \[-1y=1x-3\] after that you still want y by itself so divide -1 with everything. you should then get \[y=-x+3\] now that you have this equation you can go back to either set of inequalites. i chose the first set. So think of the y equation as a number. now you are going to have \[2x-3 (-x+3)=16\] after distributing you should have \[2x+3x-9=16\] now combine like terms and you should have \[5x-9=16\] now you add 9 to both sides and get \[5x=25\] divide 5 by both sides and x=5 now to get Y you just plug in your x (5) into one of the equations from the beginning. i chose the first one again. so you do \[2(5)-3y=16\] you then get \[10-3y=16\] subtract 10 from both sides and you should get \[3y=6\] divide three from both sides and your answer should be y=2 So x=5 and y=2

OpenStudy (anonymous):

Wow, thought it was glitched and stuck on saying you were typing I really appreciated that NB. Pg, if you would like to explain it to me I'd love to try it out

OpenStudy (anonymous):

May not be worth it though, I already gave nb the best answer. Don't want you to waste your time.

OpenStudy (anonymous):

no problem, glad i could help

OpenStudy (anonymous):

elimination is the process of "eliminating" variables from the equations in a systematic way. let's say we wanted to eliminate the x variable from the equations. we'll do this by getting the coefficients of the x terms to be opposites of each other by multiplying one or more of the equations: 2x - 3y = 16 is the same as 3(2x - 3y = 16) => 6x - 9y = 48 -3x + 2y = -19 is the same as 2(-3x + 2y = -19) => -6x +4y = -38 now we can add the equations. I know it seems a bit weird but it's all valid. 6x - 9y = 48 -6x + 4y = -38 ___________________ -5y = 10 => y = -2. We can either substitute back into one of the equations and solve for x or we can eliminate y and see what we get for x. I'll eliminate again to demonstrate the method but it's also very helpful if you get a fraction in your solution... 2x - 3y = 16 is the same as 2(2x - 3y = 16) => 4x - 6y = 32 -3x + 2y = -19 is the same as 3(-3x + 2y = -19) => -9x +6y = -57 now we can add the equations. 4x - 6y = 32 -9x + 6y = -57 ___________________ -5x = -25 => x = 5.

OpenStudy (anonymous):

so the soultion is the point (5, -2). and you can check that it satisfies both lines.

OpenStudy (anonymous):

it's all about the learning... not the points.

OpenStudy (anonymous):

Wow pgpilot had no idea about that. Would you rather use that method over substitution? Thanks for taking time to explain that.

OpenStudy (anonymous):

i like it better than substitution. you're welcome1